if t 1/2(half life period)=1/Ka(dissociation constant) what is the order of the reaction? (ans-2)


How to get it ?please help

Grade:

1 Answers

Sunil Kumar FP
9 years ago
for the reaction 2A------>P
for the second order reaction ,rate is given by =k[A]^2
Incase of second order reaction
1/At =1/A0 +kt
at t=t1/2 At =1/2A0
2/A0 =1/A0 +kt1/2
1/A0 =kt1/2
half life ,t(1/2) =1/(kA0)